OpenStudy (anonymous):

Estimate the number of gallons of gasoline consumed by the total of all automobile drivers in the U.S., per year. Suppose that there are about 3 × 108 people in the United States, approximately half of the them have cars, each car drives an average of 12,000 mi per year, and consumes a gallon of gasoline for each 20 mi

OpenStudy (anonymous):

3 x 108 is roughly 300 people. Half of them have cars. Half of 300 = 150. 150 x 12000 = 1,800,000 miles driven. Each car gets 20mpg. Solve for the # of gallons consumed.

OpenStudy (anonymous):

I meant 3 x 10^8 people sorry

OpenStudy (anonymous):

Do the same but replace 300 with 3 x 10^8 people.

OpenStudy (anonymous):

I got to that but how do I get gallons/yr

OpenStudy (anonymous):

If there are 3x10^8 people and half of them drive you'd have 1.5x10^8 people that drive. 1.5 x 10^8 cars x 12,000 miles/year = ______ total miles from all the cars in a year. Each car gets 20 miles per gallon.

OpenStudy (anonymous):

1.8 x 10^12

OpenStudy (anonymous):

\[\frac{ (1.5\times10^{8})\times 12000 }{ 20 }\]

OpenStudy (anonymous):

which is 9 x 10^10 but the answers given are 10^11 gal/yr 10^9 gal/yr 10^7 gal/yr 10^13 gal/yr

OpenStudy (anonymous):

10 x 10^11 <<<

OpenStudy (anonymous):

thank you